can you help me please, Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-3x – 5y = 61
7x – 5y = -9

Answer by mananth(16946)

-3 x -5 y = 61 .............1

7 x -5 y = -9 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
3 x 5 y = -61
7 x -5 y = -9
Add the two equations
10 x = -70
/ 10
x = -7.00
plug value of x in (1)
-3 x -5 y = 61
21 -5 y = 61
-5 y = 61 -21
-5 y = 40
y = -8.00
(-7,-8)